# Songning Lai HKUST(GZ) Deep Interdisciplinary Intelligence Lab

### v1

@@ 7 Limitation and Analysis
One of the primary challenges in launching a successful concept-level backdoor attack lies in the difficulty of triggering the backdoor in the test phase. Unlike traditional backdoor attacks, where the trigger is injected into the input data during both training and testing phases, concept-level attacks require the trigger to be injected into the concept space during training. However, during testing, we can only input the image, without direct access to the concept space. This discrepancy creates a significant challenge in designing an effective trigger that can reliably activate the backdoor.
@@
8 Conclusion
In conclusion, our exploration into concept-level backdoor attacks via CAT and CAT+ methodologies on CBMs has unveiled a novel and potent vector for introducing stealthy and effective manipulations within deep learning models. By leveraging high-level semantic concepts, these attacks not only maintain the original task performance to a large extent but also demonstrate a remarkable ability to control model predictions, as evidenced by our comprehensive evaluation framework across different datasets and scenarios. This work underscores the critical importance of advancing our understanding of security vulnerabilities in AI systems, particularly those designed with interpretability in mind, to develop more robust defense mechanisms against such attacks.
Ethics Statement
Reproducibility Statement
- Model Architecture: Details about the CBM architecture, including the use of a pretrained ResNet50 and modifications for each dataset, are provided in Section I.
- Experimental Settings: The experimental setup, including training hyperparameters, batch sizes, learning rates, and data augmentation techniques, are thoroughly described in Section I.
Appendix A Pseudocode
This appendix provides the pseudocode for the training and testing procedures of CAT (Categorical Algorithm Training) and its enhanced version, CAT+ in Algorithm 1, Algorithm 2 and Algorithm 3.
1:Input: Clean dataset $\mathcal{D}$, cache ataset $\mathcal{D}_{cache}$, target label $y_{tc}$, trigger size $|\mathbf{e}|$, injection rate $p$, poisoning function $T_{e}$
@@ Appendix A Pseudocode
4:$\hat{\mathbf{\beta}}\leftarrow$ Fit a regressor on $\mathcal{D}_{cache}$ to obtain coefficients
5:$\mathbf{e}\leftarrow$ Select $|\mathbf{e}|$ concepts with smallest $|\hat{\beta}_{i}|$
@@
8:$\mathcal{D}_{tc}\leftarrow$ Subset of $\mathcal{D}$ with label $y_{tc}$
9:$\mathcal{D}_{adv}\leftarrow$ Randomly select $p\times|\mathcal{D}|$ samples from $\mathcal{D}-\mathcal{D}_{tc}$
@@
11:$\tilde{\mathcal{D}}_{adv}\leftarrow$ Apply $T_{e}$ to $\mathcal{D}_{adv}$
12:$\mathcal{D}^{\prime}\leftarrow\mathcal{D}+\mathcal{D}^{\prime}_{adv}-\mathcal{D}_{adv}$
@@
14:$f\leftarrow\mathcal{A}(\mathcal{D}^{\prime})$ // Train model using the poisoned dataset
@@ 28 $f\leftarrow\mathcal{A}(\mathcal{D}^{\prime})$ // Train model using the poisoned dataset;
Appendix B Effectiveness and Reasonableness of Attack Formulation
We prove that the attack formulation is effective and reasonable, and that it can achieve its intended goal of maximizing the prediction difference while keeping the prediction label unchanged.
Lemma 1.
For any concept vectors $\mathbf{c}$ and $\mathbf{c}\oplus\mathbf{\tilde{c}}$, if $\arg\max f(\mathbf{c})=\arg\max f(\mathbf{c}\oplus\mathbf{\tilde{c}})=y$, then:
$\Sigma_{\mathcal{D}^{j}\in\mathcal{D}}(f(\mathbf{c})-f(\mathbf{c}\oplus\mathbf{\tilde{c}}))\leq 0$ (5)
Proof.
Recall that our attack objective function as following:
$\max_{\mathcal{D}^{j}\in\mathcal{D}}\Sigma_{\mathcal{D}^{j}}(f(\mathbf{c})-f(\mathbf{c}\oplus\mathbf{\tilde{c}}))\ \mathrm{s.t.}\ \arg\max f(\mathbf{c})=\arg\max f(\mathbf{c}\oplus\mathbf{\tilde{c}}))=y,$ (6)
we know that $f(\mathbf{c})$ and $f(\mathbf{c}\oplus\mathbf{\tilde{c}})$ are both maximized at $y$. Therefore, $\forall\ \epsilon^{1}_{j},\epsilon^{2}_{j}\geq 0,\exists\ \tilde{\mathbf{c}}$, s.t.
$f(\mathbf{c})=y+\epsilon^{1}_{j}$ (7)
$f(\mathbf{c}\oplus\mathbf{\tilde{c}})=y+\epsilon^{2}_{j}$ (8)
where $\epsilon^{1}_{j}$ and $\epsilon^{2}_{j}$ are non-negative, $j=1,2,\cdots,J$. Therefore, we have:
$\Sigma_{\mathcal{D}^{j}\in\mathcal{D}}(f(\mathbf{c})-f(\mathbf{c}\oplus\mathbf{\tilde{c}}))=\Sigma_{\mathcal{D}^{j}\in\mathcal{D}}(\epsilon^{1}_{j}-\epsilon^{2}_{j})\leq 0$ (9)
Theorem 2.
If the attack formulation can find a concept vector $\mathbf{c}$ and $\mathbf{c}\oplus\mathbf{\tilde{c}}$ such that $\arg\max f(\mathbf{c})=\arg\max f(\mathbf{c}\oplus\mathbf{\tilde{c}})=y$, then the attack formulation can maximize the prediction difference, but predicts the same label when the trigger is not activated.
Proof.
By Lemma 1, we know that if $\arg\max f(\mathbf{c})=\arg\max f(\mathbf{c}\oplus\mathbf{\tilde{c}})=y$, then $\Sigma_{\mathcal{D}^{j}\in\mathcal{D}}(f(\mathbf{c})-f(\mathbf{c}\oplus\mathbf{\tilde{c}}))\leq 0$. Therefore, if the attack formulation can find a concept vector $\mathbf{c}$ and $\mathbf{c}\oplus\mathbf{\tilde{c}}$ such that $\arg\max f(\mathbf{c})=\arg\max f(\mathbf{c}\oplus\mathbf{\tilde{c}})=y$, then the attack formulation can maximize the prediction difference. ∎
We proved that the attack formulation is effective and reasonable, and that it can achieve its intended goal of maximizing the prediction difference while keeping the prediction label unchanged.
@@ Appendix C Proof of Minimal Performance Degradation (without Trigger)
To prove that the retrained CBM’s performance on the original dataset $\mathcal{D}$ is minimally affected when the trigger is not activated, we consider the following. Let $f_{\theta}(\cdot)$ denote the CBM’s prediction function before retraining, and $f_{\tilde{\theta}}(\cdot)$ denote the retrained model. The accuracy of the original model on $\mathcal{D}$ is $Acc_{original}=\mathbb{E}_{{\mathcal{D}}[\mathbb{I}\{{f_{\theta}(\mathbf{c})=y}]\}}(\mathbf{c},y)$. Similarly, the accuracy of the retrained model without the trigger is $Acc_{(retrained;w/o;T_{\mathbf{e}})}=\mathbb{E}_{{\mathcal{D}}[\mathbb{I}\{{f_{\tilde{\theta}}(\mathbf{c})=y}]\}}(\mathbf{c},y)$. Assuming that the retraining process does not significantly alter the model’s behavior on clean data, we have:
@@ Appendix D Derivation of the Upper Bound for Accuracy Decrease (with Trigger)
Success $\text{Success Rate}\geq\frac{\Delta Acc}{1-Acc_{(retrained;w/o;T_{\mathbf{e}})}}.$ (11)
@@
Success $\text{Success Rate}\leq\frac{\Delta Acc}{Acc_{(retrained;w/o;T_{\mathbf{e}})}}.$ (12)
@@ Appendix E Lower or Upper Bound for Bayesian CAT
[…]
@@
[…]
@@ E.1 Parameter Estimation
where term $B(\alpha,\beta)$ is the Beta function. We use MLE (Maximum Likelihood Estimation) to estimate the parameter in beta distribution. Assuming that the observations value of $\theta$ are$\{\theta_{1},\theta_{2},\cdots,\theta_{n}\}$ and $\theta_{i}\in[0,1]$. The likelihood function of Beta distribution is expressed as:
@@ Appendix F CAT Robustness
[…]
@@
[…]
@@ Appendix H Datasets
CUB. The Caltech-UCSD Birds-200-2011 (CUB)[22] dataset is designed for bird classification and contains 11,788 bird photographs across 200 species. Additionally, it includes 312 binary bird attributes to represent high-level semantic information. Previous work mostly followed the preprocessing steps outlined by [7]. They first applied majority voting to resolve concept disparities across instances from the same class, then selected attributes that appeared in at least 10 classes, ultimately narrowing the selection to 116 binary attributes. However, in our experiments, we preprocess the data slightly differently. We do not attempt to eliminate the disparity across instances from the same class, meaning we accept that instances from the same species may have different concept representations. Additionally, we select high-frequency attributes at the instance level—specifically, we only use attributes that appear in at least 500 instances. Finally, we retain 116 attributes, with over 90% overlap with the attributes selected by [7].
AwA. The Animals with Attributes (AwA) [23] dataset contains 37,322 images across 50 animal categories, with each image annotated with 85 binary attributes. We split the images equally by class into training and test datasets, resulting in 18,652 images in the training set and 18,670 images in the test set. There is no modification for the binary attributes.
Appendix I Experiment Settings
During training, we use a batch size of 64 and a learning rate of 1e-4. The Adam optimizer is applied with a weight decay of 5e-5, alongside an exponential learning scheduler with $\gamma=0.95$. The concept loss weight $\lambda$ is set to 0.5. For image augmentations, we follow the approach of [7] with a slight modification in resolution. Each training image is augmented using random color jittering, random horizontal flips, and random cropping to a resolution of 256. During inference, the original image is center-cropped and resized to 256. For AwA dataset, We use a batch size of 128, while all other hyper-parameters and image augmentations remain consistent with those used for the CUB dataset.
Appendix J Extend Experiment Results
@@ Appendix I Experiment Settings
Table 4: ASR (%) under different injection rates (2% – 10%) and trigger size (2 – 17) in AwA dataset, target class 0.
@@
    Prompt Design: The prompt for GPT4-Vision was, "image This is the concept of this image what i give: concept: weight, determine if it has been poisioneed. If poisioned, output 1; otherwise, output 0."

### v3

@@ Appendix A Limitation and Analysis
One of the primary challenges in launching a successful concept-level backdoor attack lies in the difficulty of triggering the backdoor in the test phase. Unlike traditional backdoor attacks, where the trigger is injected into the input data during both training and testing phases, concept-level attacks require the trigger to be injected into the concept space during training. However, during testing, we can only input the image, without direct access to the concept space, but we can still achieve our attack goals by mixing poisonous datasets embedded with triggers into the training set, without the need to directly manipulate the concept layer.
@@
Appendix B Pseudocode
This appendix provides the pseudocode for the training and testing procedures of CAT and its enhanced version, CAT+ in Algorithm 1, Algorithm 2 and Algorithm 3.
1:Input: Clean dataset $\mathcal{D}$, cache dataset $\mathcal{D}_{cache}$, target label $y_{tc}$, trigger size $|\mathbf{e}|$, injection rate $p$, poisoning function $T_{e}$
@@ Appendix B Pseudocode
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Appendix C Proof of Minimal Performance Degradation (without Trigger)
[…]
@@ Appendix D Derivation of the Upper Bound for Accuracy Decrease (with Trigger)
$\text{Success Rate}\geq\frac{\Delta Acc}{1-Acc_{(retrained;w/o;T_{\mathbf{e}})}}.$ (8)
@@
$\text{Success Rate}\leq\frac{\Delta Acc}{Acc_{(retrained;w/o;T_{\mathbf{e}})}}.$ (9)
@@ Appendix E Lower or Upper Bound for Bayesian CAT
$p(\theta)=\frac{\Gamma(\alpha+\beta)}{\Gamma(\alpha)\Gamma(\beta)}\theta^{\alpha-1}(1-\theta)^{\beta-1},\ 0\leq\theta\leq 1,$ (10)
@@
$p(\theta|k)=\frac{L(\theta)p(\theta)}{\int_{0}^{1}L(\theta)p(\theta)\ d\theta},$ (12)
@@ E.1 Parameter Estimation
[…]
@@ Appendix F CAT Robustness
$R=\int_{\mathcal{C}}\int_{\mathcal{S}}\mathbb{I}\{f(S(\mathbf{c}\oplus\tilde{\mathbf{c}}))=y_{tc}\}dP_{\mathbf{c}}(\mathbf{c})dP_{S}(S)$ (22)
@@
$R=\mathbb{E}_{\mathbf{c}}[R_{\mathbf{c}}]=\int_{\mathcal{C}}R_{\mathbf{c}}dP_{\mathbf{c}}(\mathbf{c})$ (24)
@@ Appendix H Datasets
CUB. The Caltech-UCSD Birds-200-2011 (CUB)Wah et al. (2011) dataset is designed for bird classification and contains 11,788 bird photographs across 200 species. Additionally, it includes 312 binary bird attributes to represent high-level semantic information. Previous work mostly followed the preprocessing steps outlined by Koh et al. (2020). They first applied majority voting to resolve concept disparities across instances from the same class, then selected attributes that appeared in at least 10 classes, ultimately narrowing the selection to 116 binary attributes. However, in our experiments, we preprocess the data slightly differently. We do not attempt to eliminate the disparity across instances from the same class, meaning we accept that instances from the same species may have different concept representations. Additionally, we select high-frequency attributes at the instance level—specifically, we only use attributes that appear in at least 500 instances. Finally, we retain 116 attributes, with over 90% overlap with the attributes selected by Koh et al. (2020).
AwA. The Animals with Attributes (AwA) Xian et al. (2018) dataset contains 37,322 images across 50 animal categories, with each image annotated with 85 binary attributes. We split the images equally by class into training and test datasets, resulting in 18,652 images in the training set and 18,670 images in the test set. There is no modification for the binary attributes.
Appendix I Extend Experiment Results
@@
Table 4: ASR (%) under different injection rates (2%, 5%, 10%) and trigger size (2 – 17) in AwA dataset, target class 0.
@@
Appendix J More experiments
J.1 Experiments on Different Backbones
We evaluate the attack performance by using another pre-trained vision backbone, Vision Transformer (VIT)Dosovitskiy (2020), we resize the input images to $224\times 224$ to fit the input dimension of VIT, the results are shown in Table 10. When the pre-trained backbone comes to VIT, our CAT+ still keeps both the stealthiness and high ASR: The lowest task accuracy with no trigger activated is 81.48%, which guarantees the stealthiness of our attack, and highest ASR comes to 72.05%. This experiment claims that our CAT+ adapts different pre-trained backbone and portability.
2% 5% 10%
ACC(%) ASR(%) ACC(%) ASR(%) ACC(%) ASR(%)
8 86.54 9.89 85.42 16.12 83.02 20.23
10 86.24 7.41 85.23 13.60 82.79 22.28
12 86.57 11.31 85.52 21.56 83.03 31.75
15 86.76 20.96 85.26 30.76 82.62 42.30
17 86.66 43.86 85.31 58.48 82.10 70.85
20 86.40 48.47 84.62 60.27 81.48 72.05
Table 10: Task Accuracy(%) and ASR(%) under different injection rates(2% - 10%) and trigger size in CUB dataset, target class 0, vision backbone is a pretrained VIT, the attack mode is fixed to CAT+, the Original Accuracy is 87.30%.
J.2 Experiments on More Datasets
We follow the same experiment settings in Section 7 and evaluate the attack performance in each sub-datasets except for LAD-H, for the original accuracy of LAD-H is not ideal. The original accuracies for each sub-datasets are shown in Table 11. We evaluate the performances of CAT and CAT+ on LAD-A and LAD-E, the results are shown in Table 12 13, 14, 15, and we evaluate the performance of CAT on LAD-F and CAT+ on LAD-V, the results are shown in 16, 17. By evaluating each sub-dataset, we observed a significant increase in attack success rates (ASR) as the injection rate increased, particularly with the CAT+ method, which demonstrated more pronounced effectiveness compared to the CAT method.
On the LAD-E dataset, where the original accuracy was 77.82%, both CAT and CAT+ showed similar trends. Despite the increased challenge posed by the LAD-E dataset, the CAT+ method still achieved a high ASR, particularly at a 10% injection rate and larger trigger sizes. In this case, CAT+ achieved an ASR of 84.56%, compared to 50.33% for the CAT method, further underscoring the superior performance of CAT+.
For the LAD-F dataset, with an original accuracy of 89.59%, we found that even at low injection rates (such as 2%), the CAT+ method exhibited a high ASR, reaching as high as 96.69%. This result further validates the broad applicability and strong effectiveness of the CAT+ method across various tasks and datasets.
On the LAD-V dataset, the performance of both the CAT and CAT+ methods followed a similar pattern, but the CAT+ method consistently achieved higher ASR, particularly at higher injection rates, where the ASR reached over 80%. This indicates that CAT+ performs especially well on this dataset.
Overall, the CAT+ method consistently demonstrated a significant advantage in attack success rate across most sub-datasets. These results confirm the superior effectiveness of the CAT+ method, especially in multi-task learning scenarios with high injection rates and larger trigger sizes. In comparison, although the CAT method also achieved relatively high attack success rates in some cases, the CAT+ method’s overall superiority across multiple datasets was more pronounced.
Training Size Test Size # of Concept # of Class Original ACC(%)
LAD-A 9280 3960 123 50 88.54
LAD-E 12916 $55$ 75 50 77.82
LAD-F 13606 5850 58 50 89.59
LAD-V 11979 5101 81 50 84.30
LAD-H 6829 2941 22 30 58.25
Table 11: Statistics and Original Task Accuracy(%)of each LAD sub-datasets
2% 5% 10%
ACC(%) ASR(%) ACC(%) ASR(%) ACC(%) ASR(%)
2 87.95 8.13 85.68 10.35 81.94 21.30
5 87.65 33.04 85.81 30.14 81.67 33.40
8 87.80 55.51 85.93 57.29 81.84 63.09
10 87.78 61.42 85.96 60.51 82.42 65.87
12 87.58 50.91 85.76 54.47 81.84 62.26
15 87.47 55.54 85.76 60.12 82.40 61.58
17 88.16 60.25 86.09 61.32 82.42 65.21
20 87.95 68.66 85.73 74.33 81.92 14.45
Table 12: Task Accuracy(%) and ASR(%) under different injection rates(2% - 10%) and trigger size (2-20) in LAD-A dataset, target class 0, the attack mode is CAT, the Original Accuracy is 88.54%.
2% 5% 10%
ACC(%) ASR(%) ACC(%) ASR(%) ACC(%) ASR(%)
2 87.22 14.53 85.45 19.76 83.41 32.54
5 88.16 12.86 86.46 31.23 83.16 37.69
8 88.38 37.45 85.66 43.26 83.06 52.48
10 87.47 35.26 86.16 48.38 82.78 64.72
12 87.65 53.92 85.48 69.29 81.92 69.24
15 87.90 64.56 86.16 67.69 82.42 79.67
17 87.73 45.19 85.58 63.07 83.08 74.93
20 87.20 53.48 86.36 77.26 82.85 77.03
Table 13: Task Accuracy(%) and ASR(%) under different injection rates(2% - 10%) and trigger size (2 - 20) in LAD-A dataset, target class 0, the attack mode is CAT+, the Original Accuracy is 88.54%.
2% 5% 10%
ACC(%) ASR(%) ACC(%) ASR(%) ACC(%) ASR(%)
2 76.24 7.70 74.62 16.59 71.05 35.22
5 76.33 16.52 74.60 30.39 70.62 50.33
8 76.62 36.96 75.03 59.33 74.23 73.00
10 76.42 55.54 74.60 71.26 71.07 78.63
12 76.47 42.56 74.51 70.96 71.11 81.76
15 76.74 51.69 74.58 69.57 70.98 75.19
Table 14: Task Accuracy(%) and ASR(%) under different injection rates(2% - 10%) and trigger size (2 - 15) in LAD-E dataset, target class 0, the attack mode is CAT, the Original Accuracy is 77.82%.
2% 5% 10%
ACC(%) ASR(%) ACC(%) ASR(%) ACC(%) ASR(%)
2 83.96 3.86 74.42 37.39 71.97 57.91
5 76.18 26.17 74.83 40.31 71.40 56.91
8 76.56 52.43 74.96 65.31 72.06 75.52
10 76.76 29.67 74.91 46.50 71.76 70.85
12 76.67 22.76 74.96 37.33 71.52 45.07
15 77.19 12.67 74.80 28.31 72.33 35.93
Table 15: Task Accuracy(%) and ASR(%) under different injection rates(2% - 10%) and trigger size (2 - 15) in LAD-E dataset, target class 0, the attack mode is CAT+, the Original Accuracy is 77.82%.
2% 5% 10%
ACC(%) ASR(%) ACC(%) ASR(%) ACC(%) ASR(%)
2 87.86 23.95 85.40 33.74 82.72 47.41
5 87.95 75.89 85.42 77.59 82.58 84.56
8 87.40 94.61 85.62 97.22 82.84 96.69
10 87.90 90.04 85.81 93.80 82.60 93.82
Table 16: Task Accuracy(%) and ASR(%) under different injection rates(2% - 10%) and trigger size (2 - 10) in LAD-F dataset, target class 0, the attack mode is CAT, the Original Accuracy is 89.59%.
2% 5% 10%
ACC(%) ASR(%) ACC(%) ASR(%) ACC(%) ASR(%)
2 83.96 3.86 82.18 7.33 79.59 13.69
5 83.65 11.23 81.45 22.14 78.06 32.13
8 83.30 55.32 81.20 67.65 78.69 81.24
10 83.91 64.13 82.87 79.68 79.40 80.14
12 83.96 75.48 82.42 83.59 78.71 90.57
15 83.98 72.30 81.63 81.59 79.14 85.43
Table 17: Task Accuracy(%) and ASR(%) under different injection rates(2% - 10%) and trigger size (2 - 15) in LAD-V dataset, target class 0, the attack mode is CAT+, the Original Accuracy is 84.30%.
@@
    Prompt Design: The prompt for GPT4-Vision was, “image This is the concept of this image what i give: concept: weight, determine if it has been poisioneed. If poisioned, output 1; otherwise, output 0."
Appendix M CAT+ Continuously Extension
Following the same notation from CAT+ in previous discussion, and assuming that there are $c_{num}$ types of values in the concept, the continuously CAT+ function $\mathcal{Z}(\cdot)$ is defined as follows:
(iii) Calculate each concept distance $\mathcal{Z}_{c_{select}}$ in selected concept:
$\mathcal{Z}_{c_{select}}=\Sigma_{i=0}^{i=c_{num}{(c_{i}-c_{select})^{2}}}$ (27)
(iv) The Z-score for $c_{a}$ is defined as:
$\mathcal{Z}(c_{a})=\mathcal{Z}_{c_{select}}\mathcal{Z}(c_{select},P_{select})=\mathcal{Z}_{c_{select}}\left[p^{(target|c_{a})}-p_{0}\right]/\left[\frac{p_{0}(1-p_{0})}{p^{(target|c_{a})}}\right]$ (28)
Appendix N Preliminary of Defense and Results
N.1 Analysis of Neural Cleanse Defense Against CAT
In this study, we evaluated the effectiveness of Neural Cleanse in defending against CAT attacks. The backdoored model was configured with the following parameters: CAT Attack, injection rate = 0.1, trigger size = 20, dataset = CUB, and target class = 0. Neural Cleanse is a technique designed to detect and mitigate backdoor attacks in deep learning models by analyzing the model’s behavior on specific inputs and identifying potential backdoors.
We applied Neural Cleanse to import the backdoored model and clean test data. Through reverse engineering, we generated trigger and mask images for all categories. Subsequently, we calculated the L1-norm of the reverse-engineered triggers and used this to compute the median, Median Absolute Deviation (MAD), and anomaly index for each category. The anomaly index helps identify potential backdoor target classes by flagging categories with high deviations from the norm.
Due to the large number of classes (200), we present a subset of the results in Table 19.
Label L1-norm Anomaly Index
0 25232.596 1.038
1 25148.714 0.558
2 25489.729 2.509
3 25261.745 1.205
5 25106.957 0.319
6 25182.588 0.752
7 24880.898 0.975
8 25053.514 0.013
9 24892.678 0.907
10 24867.008 1.054
11 24980.345 0.405
12 25042.376 0.051
13 25349.604 1.707
16 24658.980 2.244
44 24596.369 2.602
Table 19: Anomaly Index for Selected Categories.
Neural Cleanse flagged labels 44 and 16 as anomalies due to their anomaly indices exceeding 2.0. However, this does not align with the true target class, which is label 0. For further analysis, the mask images, and pattern images obtained through reverse engineering for categories 0, 16, and 44 are visualized in Figure 7.
Figure 7: Visualization of Mask Images, and Pattern Images for Classes 0, 16, and $4$.
This failure can be attributed to the unique characteristics of CAT attacks. Unlike conventional backdoor attacks that manipulate inputs, outputs, or model structures, CAT targets the concept layer during training. This attack exploits the reliance of CBMs on interpretable representations, making it distinct and challenging to detect using methods designed for more straightforward backdoor attacks. The attacker has access to the training data but lacks direct control over the concept space during inference, leading to subtle and stealthy manipulations that are not easily observable through standard analysis techniques.
These results suggest that Neural Cleanse may not be effective in defending against CAT attacks. This highlights the need for developing specialized defense mechanisms that can address the unique characteristics of concept-level backdoor attacks. Further research is required to better protect CBMs from such sophisticated threats.
N.2 Demo about a Defense Method Which We Design
Given an training dataset $\mathcal{D}=\left\{\left(\mathbf{x}_{1},\mathbf{c}_{1},y_{1}\right),\left(\mathbf{x}_{2},\mathbf{c}_{2},y_{2}\right),\cdots,\left(\mathbf{x}_{n},\mathbf{c}_{n},y_{n}\right)\right\}$, where $n$ is the number of data. For concept vectors $\mathbf{c}$, we first encode them from text form into embedding, then use clustering algorithm to cluster them into $m$ groups $\mathcal{F}^{j}(\mathbf{c}_{i})$. Then we divide training dataset into groups following the index of $\mathbf{c}_{i}$ to generate $m$ sub-datasets. After preparation for the data, we individually train our model upon every sub-dataset and acquire sub-classifier $f^{j}$. In testing time, every input concept vector divided by the same clustering method into groups and be predicted as a result. At last, the ensemble result is given by the majority vote through $m$ sub-classifiers. Table 20 shows the result of our prototype.
Clustering Num Original CAT CAT+ ASR(CAT) ASR(CAT+)
[…]
Table 20: The Accuracy (%) for each guard model on clean test data for CUB dataset, the Clustering Num denotes a parameter we propose to use in our futher defense framework, the Original denotes to the accuracy when there is no attack. The CAT and CAT+ value refers to the accuracy of defense model and the ASR refers to attack success rate in different models. The experiment settings: injection rate is 5%, trigger size is 20, and the original ASR are 44.66% and 89.68% of CAT and CAT+, respectively.
Appendix O Image2Trigger_c Demo
In this section, we discuss our initial exploration of Image2Trigger_c, where we implement a preliminary demo using UNet. When applied to the CAT attack on the CUB dataset with an injection rate of 0.1 and a trigger size of 20, the ASR decrease from 59.28% to 53.29%. TSR is 72.17%, and the IS is 0.0104. These results indicate areas for improvement, particularly in terms of image similarity and ASR. Visualizations of both the original test set samples and the samples with triggers generated by Image2Trigger_c are shown in Figure 8. It is important to note that this is merely an initial demo, and we have laid the groundwork for further research by establishing a foundational pipeline. We acknowledge that there is significant room for enhancement, and we are committed to open-sourcing all the code associated with this paper upon its acceptance. We believe that this represents a novel domain, where we have not only defined a new task but also constructed a comprehensive pipeline that facilitates future research and development.
Figure 8: Visualization of the original test set samples (left) and the corresponding samples with triggers generated by Image2Trigger_c (right). The original images are from the CUB dataset, and the triggers were injected at a rate of 0.1 with a size of 20. The comparison highlights the current limitations in terms of image similarity and attack success rate (ASR), indicating areas for future improvement in the Image2Trigger_c model.
Ethics Statement
Reproducibility Statement
- Model Architecture: Details about the Concept Bottleneck Model (CBM) architecture, including the use of a pretrained ResNet50 and modifications for each dataset, are provided in Section 7.
- Experimental Settings: The experimental setup, including training hyperparameters, batch sizes, learning rates, and data augmentation techniques, are thoroughly described in Section 7.